The circumference of a circle is 31.4 cm. Find the radius and the area of the circle? (Take π = 3.14)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that
Circumference of a circle = 31.4 cm
We know that,
Circumference of a circle = 2πr
31.4 = 2 × 3.14 × r
31.4 = 6.28 × r
31.4/6.28 = r
r = 5 cm
Then,
Area of the circle =πr^2
= 3.14 × 5^2
= 3. 14 × 25
= 78.5 cm

Thanks